# LDM: Large Tensorial SDF Model for Textured Mesh Generation

### v2

@@ 1 Introduction
SDF representation has been proven capable of reconstructing high-quality. However, introducing SDF to the learning based 3D generation pipeline encounters the challenges of high memory computation and unstable convergence. Zero12345 and Zero12345++  [13, 12] introduce 3D CNNs or 3D diffusion to predict SDF volumes from image inputs. However, the quality of the output is constrained by the resolution of the volume grid, and both 3D diffusion and 3D CNNs incur significant computational overhead. As observed in CRM  [33], predicting the correct geometry using pure RGB images is extremely hard, and it adds extra geometry information into the inputs of the pipeline to help the convergence of geometry generation with the SDF representation.
@@
To generate 3D assets with high-quality meshes and illumination-decoupled textures within the end-to-end learning based framework, we propose LDM, a novel 3D generation pipeline with tensorial SDF representation and decoupled color field. Specifically, to improve the object surface quality and reduce memory requirements, we combine an SDF representation for the geometry with a low-rank tensorial representation in TensoRF[2]. To predict a decoupled tensorial representation, SDF values, albedo color, and shading color at any point in space are decoded from the tensorial representation and supervised by reference albedo and color images using volume rendering. To achieve finer texture and geometry details, a differentiable iso-surface extraction layer like Flexicube [26] can be integrated to enable high-resolution image supervision following the tensorial SDF representation.
However, the designed pipeline fails to converge due to facts as follows. 1) When converting SDF to density field as works like VolSDF [37], parameters like $\beta$ controlling the sharpness of conversion are required to adjust to current convergence status of SDF: a large $\beta$ results in the converted density field too smooth, causing the loss of geometric details; conversely, a small $\beta$ leads to geometric fragmentation. To address this issue, we introduce an adaptive $\beta$ adjusting schedule, making $\beta$ an optimizable parameter to allow $\beta$ to be adjusted directly by the gradient descent from the final losses. 2) Directly training our pipeline using the Flexicube render layer fails to converge due to the sparse view supervision of each object and gradients descent affecting only the vicinity of the SDF grid vertices. To enable the convergence, we leverage volume rendering which samples points in the entire space to stabilize the training and provide a initial SDF for the Flexicube render to refine local quality.
@@
  • We propose the first feed-forward framework capable of generating high-quality triangular meshes with illumination-decoupled RGB textures from text or a single image input in just a few seconds.
@@
Figure 1: Given a text prompt or a single image, our framework can generate corresponding high-quality 3D assets within seconds, including illumination-decoupled texture maps, facilitating integration into various downstream applications.
@@ 2 Related Work
Currently, there are some feed-forward methods capable of generating 3D assets directly from text and image inputs without optimization. LRM [5] introduced the first end-to-end 3D Large Reconstruction Model, predicting a NeRF of the object from a single image in just 5 seconds. Since LRM only uses a single image as the condition for the transformer, the resulting object’s back side exhibits sparse coloring, inconsistent with the front. Instant3D [10] improves LRM by generating multiple viewpoints from a single image through multi-view diffusion. The encoded views act as constraints to predict the object’s NeRF, enhancing the quality of 3D assets from different perspectives. 3DGS [8] has emerged as a new attempt at object representation because it can quickly produce high-quality, high-resolution images from any viewpoint through splatting and rasterization. LGM [30]encodes attributes of 3D Gaussians from multiple views into splatter images [29] and uses a UNet to predict these images from RGB inputs.TMGS [40] uses a transformer to predict tri-plane attribute fields, enriching point cloud models of objects with additional properties to achieve a 3DGS representation. Nevertheless, both 3DGS and NeRF representations encounter difficulties with issues of uneven geometries and suboptimal geometric quality. SDF representation has been proven capable of reconstructing high-quality geometries Zero12345 and Zero12345++  [13, 12] use 3D CNNs or 3D diffusion to predict SDF volumes from images, but the output quality is limited by the volume grid’s resolution. CRM [33] uses diffusion to predict six Canonical Coordinates Maps from a single input and optimizes a UNet to predict a tri-plane from these CCMS, which includes an SDF field. CRM integrates a differentiable iso-surface extraction layer [26] to optimize the generated SDF. These efforts have produced impressive textured meshes; however, their textures embed shading effects such as shadows and highlights, making these meshes unsuitable for relighting and material editing. This limitation hinders their use in downstream applications.
@@ 3.2 Tensorial SDF and Decoupled Color Field
$V_{p}=V^{X}_{k}\circ M^{YZ}_{k}\oplus V^{Y}_{k}\circ M^{ZX}_{k}\oplus V^{Z}_{k}\circ M^{XY}_{k}$ (1)
where $V_{p}$ denoted the feature vector of position $p$. In addition, $V^{m}_{k}$ and ${M}_{k}^{\tilde{m}}$ represent the $k$-th vector and matrix factors of their corresponding spatial axes $m$, and $\tilde{m}$ is the two axes orthogonal to m (e.g., $\tilde{X}=YZ$). $\circ$ and $\oplus$ denote the element-wise multiplication and concatenation operations. Moreover, unlike TensoRF, which uses two separate tensor fields to encode geometry and appearance, we leverage a single shared tensorial field for both. This approach enhances the correlation between geometry and appearance.
Model SDF and decoupled color field. We aim to decompose the final rendering color into two parts. The first part is the albedo color, which reflects the inherent color of the object and remains constant for each object. The second part is the shading color, which is determined by the varying lighting conditions the object receives and its different material properties. Given the continuous nature of the decomposed appearance field and SDF field, we introduce multiple MLP decoders to decode the purely explicit features extracted from the tensorial representation. This approach not only prevents the unsmoothness that might arise from purely explicit representations but also allows for the prediction of various material properties and SDF from compact features. Formally,
@@
where $\Theta_{s},\Theta_{c}$ where represent the MLPs used to decode the SDF value, albedo color, and shading color. Equipped with the albedo color $c_{a}$ and shading colors $c_{s}$, the final rendering color $c$ of point $p$ can be calculated as $c=c_{a}\cdot c_{r}$. Once the albedo map with decoupled illumination is obtained, the object can be used in various downstream applications. In new scenes, the shading color can be calculated based on different materials and lighting conditions, following the rendering equation [7].
@@ 3.3 Feed-forward Large Reconstruction Model
Specifically, we use a vision transformer, DINO2 [1], to encode feature tokens $F_{I}$ from multi-view images as shown in Figure 2. To ensure that the resulting feature tokens incorporate multi-view camera information, we modulate the camera information as described in LRM [5] and inject it into the image encoder using AdaLN [6, 21]. Furthermore, we arrange the tensor vectors $V^{m}_{k}$ and ${M}_{k}^{\tilde{m}}$ into learnable tokens and feed these tokens into a feature decoder composed of a sequence of transformer layers, resulting in a series of tensor tokens. It should be noted that the information from the conditioned image feature $F_{I}$ is connected to the cross-attention layer within these transformer layers. Finally, we reshape and upsample the tensor tokens into the final tensorial SDF representation. Then, all the transformer architecture can be trained in an end-to-end manner using image reconstruction loss at novel views images rendered from the tensorial SDF representation using volume rendering [19]. More details of the network structure can be found in the appendix. The training loss in this stage is defined as,
$\mathcal{L}_{1}$ $=\mathcal{L}_{mse}(I_{rgb},I^{GT}_{rgb})+\mathcal{L}_{mse}(I_{\alpha},I^{GT}_{\alpha})+\mathcal{L}_{mse}(I_{h},I^{GT}_{h})+\lambda_{vgg}\mathcal{L}_{lpips}(I_{rgb},I^{GT}_{rgb})$ (4)
where $I_{rgb},I_{\alpha},I_{\alpha}$ respectively represent the predicted final color image, albedo image, and mask image. Following LRM [5], we apply mean square error loss and VGG-based LPIPS loss [38] to the predicted images and corresponding reference images. During training, we set $\lambda_{\text{vgg}}=2$.
@@ 3.4 Lifting SDF to Fine Mesh
Till now, we have introduced a transformer structure that enables the generation of a tensorial SDF representation from sparse images. We can directly extract geometric surfaces from the SDF representation using Marching Cubes [18] after training. However, the strategy has the issues: 1) given the high computational cost of volume rendering, only a patch of the image is rendered during training, thus being unable to fully utilize full-resolution supervisory images. 2) a gradient-based mesh optimization layer after a SDF field can improve the silhouette quality as described in Nvdiffrec [20].
@@
$\mathcal{L}_{2}$ $=\mathcal{L}_{1}+\lambda_{d}\left\|I_{d}-I_{d}^{gt}\right\|_{1}+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}}$ (6)
@@ 4.1 Implementation Details
4.2 Comparison
We compare our method with the previous state-of-the-art feed-forward generation methods, including One-2-3-45 [14], LRM [5], CRM [33] and LGM [30]. We evaluate our method on the Google Scanned Object (GSO) dataset [4]. To quantitatively evaluate the image quality synthesized by our approach, we adopt three standard metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [34], and Learned Perceptual Image Patch Similarity (LPIPS) [38]. Regarding geometric quality, we report the Chamfer Distance (CD) and Volume Intersection over Union (IoU) as metrics. The quantitative results, as shown in Table 2, demonstrate that our method outperforms others in both color and geometry. Additionally, we show a qualitative comparison of various methods in the Figure 4. As shown in the figure, these methods have all achieved good results, but some of them struggle with certain detailed aspects. One2345 [14] is limited by the discrete SDF volume grid representation, resulting in blurred textures and geometry. LRM [5], lacking multi-view images as constraints, produces 3D models with inconsistent back and front colors and suffers from multi-face Janus problem [28]. LGM [30], using 3DGS as its representation, tends to produce blurred details in areas with dense geometry, such as leaf clusters. CRM [33] exhibits geometric discontinuities in the frog’s hands and the flower’s petioles. However, our method successfully generates 3D assets that are well-aligned with the input conditions, featuring smooth and intact geometric structures.
@@ 4.3 Ablation Study
Different object representation.
@@
Flexicubes layer
We conducted ablation experiments to verify the impact of introducing the Flexicube layer for model fine-tuning. As shown in  Figure 9 and  Table 7, training with the Flexicubes layer effectively enhances the texture clarity, thanks to the utilization of higher resolution during training. In addition, introducing depth constraints leads to smoother geometric surfaces. For more ablation discussions, please refer to the appendix  Section 6.2.
@@
Figure 7: Quantitative effect of the Flexlayer, evaluate the performance of the fine-tuned model with varying numbers of input views. More discussion can be find in  Section 6.2.

### v3

@@ Abstract
Keywords: 3D Generation model, Diffusion, Intrinsic decomposition, Relighting.
@@ 1 Introduction
Figure 1: Given a text prompt or a single image, our framework can generate corresponding high-quality 3D assets within seconds, including illumination-decoupled texture maps, facilitating integration into various applications, such as relighting and material editing.
@@
SDF representation has been proven capable of reconstructing high-quality. However, introducing SDF to the learning based 3D generation pipeline encounters the challenges of high memory computation and unstable convergence. Zero12345 and Zero12345++  [16, 15] introduce 3D CNNs or 3D diffusion to predict SDF volumes from image inputs. Nevertheless, the quality of the output is constrained by the resolution of the volume grid, and both 3D diffusion and 3D CNNs incur significant computational overhead. As observed in CRM  [37], predicting the correct geometry using pure RGB images is extremely hard, and it adds extra geometry information into the inputs of the pipeline to help the convergence of geometry generation with the SDF representation.
@@
To generate 3D assets with high-quality meshes and illumination-decoupled textures within the end-to-end learning based framework, we propose LDM, a novel 3D generation pipeline with tensorial SDF representation and decoupled color field, as shown in Figure 1. Specifically, to improve the object surface quality and reduce memory requirements, we combine an SDF representation for the geometry with a low-rank tensorial representation in TensoRF[3]. Furthermore, to adapt to the light decoupling task in the generative approach, we introduce a shading model that separates the final output into an albedo color and a shading color.
To predict a decoupled representation, SDF values, albedo color, and shading color at any point in space are decoded from the tensorial representation and supervised by reference albedo and color images using volume rendering. However, the method used to convert SDF to a density field in reconstruction works, such as VolSDF [41], is not applicable to our generative approach. In these works, a parameter $\beta$ is required to control the sharpness of the conversion and must be adjusted based on the convergence status of the SDF for each scene, while a large $\beta$ means smoother geometry. Unlike reconstruction tasks that focus on a single object or scene, the generative approach needs to handle a variety of geometries, some smooth and others highly detailed. To address this issue, we introduce an adaptive $\beta$ adjusting schedule, making $\beta$ an optimizable parameter to allow $\beta$ to be adjusted directly by the gradient descent from the final losses.
To achieve finer texture and geometry details, a differentiable iso-surface extraction layer like Flexicube [29] can be integrated to enable high-resolution image supervision following the tensorial SDF representation. Specifically, directly training to predict the SDF field using the Flexicube [29] rendering layer fails to converge due to sparse view supervision for each object and gradient descent impacting only the vicinity of the SDF grid vertices. To address this, a preliminary volume rendering training phase is essential. This phase samples points throughout the entire space to stabilize training and provide an initial SDF, allowing the Flexicube renderer to refine local quality. Extensive experiments show that our method is capable of generating diverse, high-quality 3D mesh assets along with decomposed RGB textures in just seconds.
@@
  • We propose the first feed-forward framework capable of generating high-quality meshes with illumination-decoupled RGB textures from text or a single image input in just a few seconds.
@@ 2 Related Work
Currently, there are some feed-forward methods capable of generating 3D assets directly from text and image inputs without optimization. LRM [6] introduced the first end-to-end 3D Large Reconstruction Model, predicting a NeRF of the object from a single image in just 5 seconds. Since LRM only uses a single image as the condition for the transformer, the resulting object’s back side exhibits sparse coloring, inconsistent with the front. Instant3D [11] improves LRM by generating multiple viewpoints from a single image through multi-view diffusion. The encoded views act as constraints to predict the object’s NeRF, enhancing the quality of 3D assets from different perspectives. 3DGS [10] has emerged as a new attempt at object representation because it can quickly produce high-quality, high-resolution images from any viewpoint through splatting and rasterization. LGM [33]encodes attributes of 3D Gaussians from multiple views into splatter images [32] and uses a UNet to predict these images from RGB inputs.TMGS [44] uses a transformer to predict tri-plane attribute fields, enriching point cloud models of objects with additional properties to achieve a 3DGS representation. Nevertheless, both 3DGS and NeRF representations encounter difficulties with issues of uneven geometries and suboptimal geometric quality. SDF representation has been proven capable of reconstructing high-quality geometries Zero12345 and Zero12345++  [16, 15] use 3D CNNs or 3D diffusion to predict SDF volumes from images, but the output quality is limited by the volume grid’s resolution. CRM [37] uses diffusion to predict six Canonical Coordinates Maps from a single input and optimizes a UNet to predict a tri-plane from these CCMS, which includes an SDF field. CRM integrates a differentiable iso-surface extraction layer [29] to optimize the generated SDF. Concurrent work InstantMesh [38] utilizes an LRM-based model to predict 3D assets represented as NeRF from multi-view images produced by multi-view diffusion. These NeRFs are subsequently refined into a Signed Distance Field (SDF). These efforts have produced impressive textured meshes; however, their textures embed shading effects such as shadows and highlights, making these meshes unsuitable for relighting and material editing. This limitation hinders their use in downstream applications.
@@ 3.2 Tensorial SDF and Decoupled Color Field
$V_{p}=V^{X}_{k}\circ M^{YZ}_{k}\oplus V^{Y}_{k}\circ M^{ZX}_{k}\oplus V^{Z}_{k}\circ M^{XY}_{k},$ (1)
where $V_{p}$ denoted the feature vector of position $p$. In addition, $V^{m}_{k}$ and ${M}_{k}^{\tilde{m}}$ represent the $k$-th vector and matrix factors of their corresponding spatial axes $m$, and $\tilde{m}$ is the two axes orthogonal to m (e.g., $\tilde{X}=YZ$). $\circ$ and $\oplus$ denote the element-wise multiplication and concatenation operations. Moreover, unlike TensoRF, which uses two separate tensor fields to encode geometry and appearance, we leverage a single shared tensorial field for both. This approach enhances the correlation between geometry and appearance, as mentioned in [12].
Model SDF and decoupled color field. Our shading model differs from traditional PBR-based shading formulations and is more akin to intrinsic image decomposition works [13, 1, 8], where the color is divided into a reflectance component (albedo) and a shading component that captures illumination information in image space. This approximation has been proven to be a fast and effective approach. Using PBR-based shading models requires modeling and predicting the ambient lighting of the input condition images when generating 3D assets, which could lead to unstable model training and difficulty in convergence. Additionally, the condition images are generated by a diffusion model, they cannot ensure consistent lighting across multiple images, which further complicates the direct prediction of lighting. Adopting a simplified shading method that is closer to intrinsic image decomposition is a trade-off approach for use in feed-forward generative models.
Therefore, we decompose the final rendering color into two parts. The first part is the albedo color, which reflects the inherent color of the object and remains constant for each object. The second part is the shading color, which is determined by the varying lighting conditions the object receives and its different material properties. Given the decomposed appearance field and SDF field, we introduce multiple MLP decoders to decode the features extracted from the tensorial representation. This approach enables predicting various material properties and SDF from compact features. Formally,
@@
where $\Theta_{s},\Theta_{c}$ represent the MLPs used to decode the SDF value, albedo color, and shading color. Equipped with the albedo color $c_{a}$ and shading colors $c_{s}$, the final rendering color $c$ of point $p$ can be calculated as,
$c=c_{a}\cdot c_{r}.$ (3)
Once the albedo map with decoupled illumination is obtained, artists can enhance the 3D assets with PBR materials such as metallic and roughness, tailoring them to specific use cases and making them ready for application. In new scenes, the shading color can be calculated based on different materials and lighting conditions, following the rendering equation [9].
@@ 3.3 Feed-forward Large Reconstruction Model
Specifically, we use a vision transformer, DINO2 [2], to encode feature tokens $F_{I}$ from multi-view images as shown in Figure 2. To ensure that the resulting feature tokens incorporate multi-view camera information, we modulate the camera information as described in LRM [6] and inject it into the image encoder using AdaLN [7, 24]. Furthermore, we arrange the tensor vectors $V^{m}_{k}$ and ${M}_{k}^{\tilde{m}}$ into learnable tokens and feed these tokens into a feature decoder composed of a sequence of transformer layers, resulting in a series of tensor tokens. It should be noted that the information from the conditioned image feature $F_{I}$ is connected to the cross-attention layer within these transformer layers. Finally, we reshape and upsample the tensor tokens into the final tensorial SDF representation. Then, all the transformer architecture can be trained in an end-to-end manner using image reconstruction loss at novel views images rendered from the tensorial SDF representation using volume rendering [22]. More details of the network structure can be found in the supplementary material. The training loss in this stage is defined as,
$\mathcal{L}_{1}$ $=\mathcal{L}_{mse}(I_{rgb},I^{GT}_{rgb})+\mathcal{L}_{mse}(I_{\alpha},I^{GT}_{\alpha})$ (5)
$+\mathcal{L}_{mse}(I_{h},I^{GT}_{h})+\lambda_{vgg}\mathcal{L}_{lpips}(I_{rgb},I^{GT}_{rgb}),$
where $I_{rgb},I_{\alpha},I_{h}$ respectively represent the predicted final color image, albedo image, and mask image. Following LRM [6], we apply mean square error loss and VGG-based LPIPS loss [42] to the predicted images and corresponding reference images. During training, we set $\lambda_{\text{vgg}}=2$.
@@ 3.4 Lifting SDF to Fine Mesh
Till now, we have introduced a transformer structure that enables the generation of a tensorial SDF representation from sparse images. We can directly extract geometric surfaces from the SDF representation using Marching Cubes [21] after training. However, the strategy has the issue that given the high computational cost of volume rendering, only a patch of the image is rendered during training, thus being unable to utilize full-resolution supervisory images fully. In addition, a gradient-based mesh optimization layer after an SDF field can improve the silhouette quality as described in Nvdiffrec [23].
@@
$\mathcal{L}_{2}$ $=\mathcal{L}_{1}+\lambda_{d}\left\|I_{d}-I_{d}^{gt}\right\|_{1}+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}},$ (7)
@@ 4 Experiment
Figure 5: The effect of illumination decoupled texture. We perform relighting in new scenes for both illumination-decomposed textures and non-decomposed textures. The 3D assets without illumination decomposition display incorrect shadows in the new scenes.
@@ 4.1 Implementation Details
Training details
We train our model in two stages. In the first stage, we train using a volume rendering pipeline and loss $\mathcal{L}_{1}$ defined as Equation (5). Specifically, we render random patches in size 128×128 cropped from the original 512×512 resolution images during training, which conserves GPU memory while increasing local resolution. However, we observed that initiating training with too large a scaling factor can impede model convergence. Therefore, we progressively increase the original resolution linearly with epochs, from 192 to 512. Additionally, we increase the likelihood of selecting crops covered by the foreground mask. This stage takes 4 days with batch size 32. We used the Adam optimizer with a learning rate of 4e-4, weight decay of 0.05, and betas of (0.9, 0.95). The learning rate is cosine annealed to 0 during the training. In the second stage, we train the model with Flexicube pipeline with full resolution using the loss defined as Equation (7). This stage takes 2 days with batch size 32. We use the same optimization settings as in the first stage, merely adjusting the learning rate to 1e-5. All training is conducted on 16 NVIDIA A6000 48GB GPUs.
4.2 Comparison
We compare our method with the previous state-of-the-art feed-forward generation methods, including One-2-3-45 [17], LRM [6], CRM [37] and LGM [33]. We compare the different features of various methods in Table 1. We evaluate our method on the Google Scanned Object (GSO) dataset [5]. To quantitatively evaluate the image quality synthesized by our approach, we adopt three standard metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [36], and Learned Perceptual Image Patch Similarity (LPIPS) [42]. Regarding geometric quality, we report the Chamfer Distance (CD) and Volume Intersection over Union (IoU) as metrics. The quantitative results, as shown in Table 2, demonstrate that our method outperforms others in both color and geometry. Additionally, we show a qualitative comparison of various methods in the Figure 4. As shown in the figure, these methods have all achieved good results, but some of them struggle with certain detailed aspects. One2345 [17] is limited by the discrete SDF volume grid representation, resulting in blurred textures and geometry. LRM [6], lacking multi-view images as constraints, produces 3D models with inconsistent back and front colors and suffers from multi-face Janus problem [31]. LGM [33], using 3DGS as its representation, tends to produce blurred details in areas with dense geometry, such as leaf clusters. CRM [37] exhibits geometric discontinuities in the frog’s hands and the flower’s petioles. However, our method successfully generates 3D assets that are well-aligned with the input conditions, featuring smooth and intact geometric structures.
@@ 4.3 Ablation Study
The effect of illumination decoupled texture.
As shown in Figure 5, We generate 3d asserts using LDM, applying both illumination-decoupled albedo texture and original RGB texture. Then, we relight the generated results in new scenes. It can be observed that the area highlighted by the red arrow contains excessive shadows, which is incorrect. Without proper illumination decomposition, the shadows baked into the RGB texture from the original lighting combine with the shadows of the new scene, leading to inaccurate results. Therefore, the generation of illumination-decoupled 3D assets plays a crucial role in the usability of downstream applications.
Comparisons across different object representation.
@@
The effect of Flexicubes layer
We conducted ablation experiments to verify the impact of introducing the Flexicube layer for model fine-tuning. As shown in  Figure 8 and  Table 3, training with the Flexicubes layer effectively enhances the texture clarity, thanks to the utilization of higher resolution during training. In addition, introducing depth constraints leads to smoother geometric surfaces.
Figure 8: Ablation study shows that the Flexicubes layer helps improve the texture details of generated objects.
Table 3: Quantitative effect of the Flexlayer, evaluate the performance of the fine-tuned model with varying numbers of input views. More discussion can be find in  Section 4.3.
@@
Figure 9: Relighting results of 3d assets generated by our framework.
Figure 10: blation study for transferring the model to a different number of inputs views shows that good results can be achieved with simple fine-tuning.
Comparisons across different diffusion methods
As most text and image-conditioned multi-view diffusion models generate 4 different views, our framework is trained under the assumption of having 4 input views. We evaluated the performance of our framework when switching between multi-view images generated by different diffusion models. Specifically, we try to switch the input of our model to the image predicted by popular multi-view diffusion models Zero123plus [30]. The experimental results, as shown in Figure 10 and  Table 3, indicate that when we directly employ a model trained with 4 views to reconstruct 6 views as input, it does not perform well and produces some misaligned results. However, after fine-tuning our model with 6 views input for only 10k iterations, it is able to generate reliable results. This experiment demonstrates that our framework exhibits good generalization across different multi-view diffusion results.
Figure 11: We applied different material textures to the generated 3D assets and modified the albedo color to produce some interesting results, such as a metallic hamburger.
5 Application
As our method generates 3D assets with illumination-decomposed texture maps, it easily supports applications such as relighting and material editing. As shown in Figure 9, we relight the 3D assets in different new scenes, producing convincing synthesized images. Additionally, we apply PBR material textures to the 3D assets and edit their metallicity, roughness, and albedo properties, with the resulting images shown in Figure 11.
Figure 12: More predicted results for text-to-3D generation task.
6 Limitations
While our model can generate high-quality illumination-decoupled 3D assets, there are still some limitations. Like other transformer-based approaches [11, 6], the size of the tensorial SDF tokens produced by our model is capped at 64x64, constraining the resolution of the final 3D assets. In addition, our illumination-decoupled module is not designed to handle complex materials, such as translucent surfaces, due to our simplified rendering composition. Finally, the quality of our reconstructed 3D models is influenced by the inconsistencies in the multi-view diffusion. We reserve these challenges for future exploration. Further discussion can be found in the appendix Section C.